how to factor with factor formula :
6x^2-x-2=0
please give step by step solution.

Answer by Alan3354(69443)

how to factor with factor formula :
6x^2-x-2=0
---------------
It's an equation. Equations are usually solved, not factored.
If you want to factor the polynomial:
------
It'll be either
(x + m)*(6x + n)
or
(2x + m)*(3x + n)
--------
In either case, m & n = 1 & 2, and the signs of m & n are different.
For (x + m)*(6x + n), only n = 1 gives an odd coefficient for x, and that doesn't work, so it's (2x + m)*(3x + n)
To get an odd coefficient for x, m has to odd.
--> (2x + 1)*(3x - 2) or (2x - 1)*(3x + 2)
(2x + 1)*(3x - 2) works.
-----
(2x + 1)*(3x - 2) = 0
-------
idk what the "factor formula" is.